# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 6.1. Challenges
ML bias in ADA: ML bias in ADA is a growing source of concern (Cheong et al., 2023a, 2021), including variability in depression manifestation (Cheong et al., 2023a; Cameron et al., 2024) and the small dataset challenge (Cheong et al., 2024c). Although some existing works have made progress on this front (Cameron et al., 2024; Cheong et al., 2024b, a), there is still a need to develop principled methods to account for other sources of bias such as labeling bias (Cheong et al., 2021) and address broader notions of fairness beyond parity-based fairness measures (Barocas et al., 2023; Cheong et al., 2022; Kuzucu et al., 2024; Cheong et al., 2021). Moreover, no existing work has investigated the interpretability, explainability, nor reproducibility of such models.
@@
Lack of interpretability and reproducibility: Most advanced DL-based ADA approaches directly infer depression predictions from human behaviours. Despite their effectiveness, these ‘black-box’ models lack explicit rule-based logic explaining how they extract depression-related cues. Specifically, they fail to clarify: (i) which behavioural patterns and their combinations are most predictive of depression, (ii) the complementary cues provided by different modalities; and (iii) why certain models outperform others in identifying depression-related behaviours. Furthermore, most of them have limited code availability, making them difficult to be fully re-produced, i.e., only a few of them have fully opensourced their implementations (Song et al., 2022a; Gimeno-Gómez et al., 2024; Xu et al., 2024; Islam and Bae, 2024; Pan et al., 2024b; Shen et al., 2022).
@@
Lack of generic solutions to tackle with common data issues: Although some approaches (Song et al., 2022a; Xu et al., 2024) attempted to learn video-level fixed-size representations, DL-based ADA models typically struggle with processing long and variable-length face videos without significant information lose/distortion, due to DL models’ fixed-size input requirements and high computational costs of processing all frames. Second, there lacks standardised and solid solutions to handle noises within depression data, e.g., low signal-to-noise EEG, speech interference, or missing target faces in video frames, despite some efforts have been already made (Song et al., 2024; He et al., 2022c). Third, cultural and linguistic diversity limits the generalizability of language-based ADA models (Ye et al., 2021; Trotzek et al., 2018; Cameron et al., 2024), resulting in poor performances under different linguistic and cultural contexts.
@@
Lack of large-scale and high-quality ADA datasets: Another common challenge is the lack of large-scale, high-quality and well-annotated publicly available datasets for researchers to develop and fairly compare their multi-modal expressive behaviour-based ADA models. Depression labels are typically obtained from self-reported questionnaires rather than standardised clinical assessments which not only limit the advancement of DL-based ADA models but also undermine evaluation objectivity (please refer to Sec. 5 for more details). Alternatively, collecting a high-quality and large-scale depression dataset with diverse participant demographics and reliable depression labels remains challenging due to ethical, financial, data anonymisation, data collection resource and high-quality clinical labeling constraints.
6.2. Opportunities and future directions
@@
There are multiple avenues for enhancing DL-based ADA solutions. First, researchers could start adding a bias analysis section in their ADA publications (e.g., similarly to the recommendations made within the affective computing community (Cheong et al., 2023b)). Second, further research could focus on developing hand-crafted behavioural descriptors for depression, grounded in physiological and psychological insights. Such descriptors would not only enhance interpretability but also guide the extraction of meaningful features for depression detection. Third, researchers can develop generalizable strategies to identify which input behaviours contribute most significantly to ADA, and which behaviours do not effectively distinguish between depressed and non-depressed individuals. Fourth, researchers could explore simulating an individual’s internal, personalized cognition based on their expressive behaviours to infer depression status - an approach inspired by similar strategies used in true personality recognition tasks (Song et al., 2022b, 2021). Fifth, future work could focus on reliably eliciting depression-related behaviours in an unobtrusive manner to enable more efficient and natural data collection. Additionally, establishing a research community around the ADA problem encouraging open science would facilitate fair comparisons, reproducibility, and broader collaboration. In summary, our work aims to contribute to the ADA field by offering a comprehensive overview, identifying current limitations, and outlining opportunities to foster cohesive collaboration and promote more ethical and impactful ADA research.

### v2

@@ 6.1 Challenges
ML bias in ADA: This is a growing source of concern [299, 300], including variability in depression manifestation [299, 301] and the small dataset challenge [249]. Although some works have made progress on this front [301, 244, 302], there is still a need to develop principled methods to account for other sources of bias such as labeling bias [300] and address broader notions of fairness beyond parity-based fairness measures [303, 304, 305, 300]. Moreover, no existing work has investigated the interpretability, explainability, nor reproducibility of such models.
@@
Lack of interpretability and reproducibility: Most advanced DL-based ADA approaches end-to-end infer depression predictions from human behaviours. Despite their effectiveness, these ‘black-box’ models lack explicit rule-based logic explaining how they extract depression-related cues and make predictions. Specifically, they fail to clarify: (i) which behavioural patterns and their combinations are most predictive of depression, (ii) the complementary cues provided by different modalities; and (iii) why certain models outperform others in identifying depression-related behaviours. Furthermore, most of them have limited code availability, making them difficult to be fully re-produced, i.e., only a few of them have fully opensourced their implementations [42, 20, 28, 215, 203, 69].
@@
Lack of generic solutions to tackle with common data issues: Although some approaches [42, 28] attempted to learn video-level fixed-size representations, DL-based ADA models typically struggle with processing long and variable-length face videos without significant information lose/distortion, due to DL models’ fixed-size input requirements and high computational costs of processing all frames. Second, there lacks standardised and solid solutions to handle noises within depression data, e.g., low signal-to-noise EEG, speech interference, or missing target faces in video frames, despite some efforts have been already made [195, 255]. Third, cultural and linguistic diversity limits the generalizability of language-based ADA models [243, 235, 301], resulting in poor performances under different linguistic and cultural contexts.
@@
Lack of large-scale and high-quality ADA datasets: Another vital challenge is the lack of large-scale, high-quality and well-annotated publicly available datasets for researchers to develop and fairly compare their expressive behaviour-based ADA models. Depression labels are typically obtained from self-reported questionnaires rather than standardised clinical assessments which not only limit the advancement of DL-based ADA models but also undermine evaluation objectivity (please refer to Sec. 5 for more details). Alternatively, collecting a high-quality and large-scale depression dataset with diverse participant demographics and reliable depression labels remains challenging due to ethical, financial, data anonymisation, data collection resource and high-quality clinical labeling constraints.
6.2 Opportunities and future directions
@@
There are multiple avenues for enhancing DL-based ADA solutions. First, researchers could start adding a bias analysis section in their ADA publications (e.g., similarly to the recommendations made within the affective computing community [306]). Second, further research could focus on developing hand-crafted behavioural descriptors for depression, grounded in physiological and psychological insights. Such descriptors would not only enhance interpretability but also guide the extraction of meaningful features for depression detection. Third, researchers can develop generalizable strategies to identify which input behaviours contribute most significantly to ADA, and which behaviours do not effectively distinguish between depressed and non-depressed individuals. Fourth, researchers could explore simulating an individual’s internal, personalized cognition based on their expressive behaviours to infer depression status - an approach inspired by similar strategies used in true personality recognition tasks [307, 308]. Fifth, future work could focus on reliably eliciting depression-related behaviours in an unobtrusive manner to enable more effective and natural data collection. Finally, establishing a research community around the ADA problem encouraging open science would facilitate fair comparisons, reproducibility, and broader collaboration. In summary, our work aims to contribute to the ADA field by offering a comprehensive overview, identifying current limitations, and outlining opportunities to foster cohesive collaboration and promote more ethical and impactful ADA research.